# QuaLLM: An LLM-based Framework to Extract Quantitative Insights from Online Forums

### v1

@@ Abstract.
Online discussion forums provide crucial data to understand the concerns of a wide range of real-world communities. However, the typical qualitative and quantitative methods used to analyze those data, such as thematic analysis and topic modeling, are infeasible to scale or require significant human effort to translate outputs to human readable forms. This study introduces QuaLLM, a novel LLM-based framework to analyze and extract quantitative insights from text data on online forums. The framework consists of a novel prompting methodology and evaluation strategy. We applied this framework to analyze over one million comments from two Reddit’s rideshare worker communities, marking the largest study of its type. We uncover significant worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls about worker insights. In short, our work sets a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; May 12, 2024; Honolulu, HI, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Figure 1. Overview of QuaLLM: an LLM-Based Framework to Extract Quantitative Insights from Online Forums
@@
Text data from online forums, such as Reddit, has been used to study issues affecting diverse communities (Proferes et al., 2021). Researchers have leveraged this data across various domains, including healthcare (Zamanifard and Robb, 2023; Xu et al., 2023; Ho et al., 2017; Paxman, 2021), political discourse (Papakyriakopoulos et al., 2023; De Francisci Morales et al., 2021; Marchal, 2020), education (Madsen and Madsen, 2022; Park et al., 2018), and gig work111a market characterized by short-term and freelance work like food delivery and ridesharing (Rosenblat and Stark, 2016; Sannon et al., 2022; Yao et al., 2021; Watkins, 2022; Ma et al., 2018), to highlight critical concerns and trends.
@@
Studies analyzing this text data for social phenomena often use interpretive qualitative methods(Gichuru, 2017) like coding (Saldaña, 2021) and thematic analysis (Braun and Clarke, 2012), to extract themes. Although detailed, these methods demand considerable human effort and interpretation. Such detailed interpretations may not always be necessary for identifying broad trends across large online communities. Alternatively, positivist quantitative (Park et al., 2020) techniques like LDA (Abdul et al., 2018) and BERT derivatives for topic modeling also distill themes from text, but offer greater scalability. However, they require careful hyperparameter selection, operate on limited context length, and require significant human effort to translate outputs to human readable forms.
@@
Recent advancements in LLMs and prompt engineering (Pham et al., 2023; Nori et al., 2023; OpenAI, 2023; Deng et al., 2023) offer an alternative for analyzing extensive text data from online forums, addressing limitations of scale and context. Specifically, LLMs have been shown to perform well on generation, classification, and ranking tasks (Zhao et al., 2023), and thus have the potential to assist researchers in text data analysis. However, the hitherto absence of a robust methodology and evaluation strategy for effectively using LLMs for online text data analysis limits its use and underscores a significant gap in the research landscape.
Our work addresses this gap through the following contributions:
@@
    We introduce QuaLLM, an LLM-based framework consisting of a novel prompting methodology and evaluation strategy for the analysis and extraction of quantitative insights from text data on online forums.
@@
    We apply our framework to a case study on Reddit’s rideshare communities, analyzing over one million comments—the largest study of its kind—to identify worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls (White House (2023), OSTP).
@@
Taken together, by establishing a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums, we hope our work can significantly assist quantitative researchers across various fields.
2. Related Work
@@
LLM Prompting and Evaluation.   LLMs have been used for generation (Pham et al., 2023; Zhao et al., 2023), classification (Deng et al., 2023), and ranking (Qin et al., 2023) tasks via advances in prompt engineering (Nori et al., 2023; OpenAI, 2023). Popular prompting methodologies include: In-context learning (Brown et al., 2020), which teaches LLMs new tasks with few labeled examples, Chain of Thought (Wei et al., 2022), which breaks down complex problems through stepwise reasoning, and Ensembling, (Wang et al., 2022) which aggregates outputs of multiple model runs to obtain a consensus output. Despite these prompting advances, accurately evaluating LLM outputs is challenging, with human evaluation generally preferred (Amer, 2023). Building on prior work, we introduce a framework consisting of a novel prompting methodology and evaluation strategy for quantitative analysis of text data from online forums.
@@
Gig Work and Reddit Analysis.   Regulatory bodies (White House (2023), OSTP) have sought public insights on the impact of AI and algorithmic management by employers (Kantor and Sundaram, 2022; Bernhardt et al., 2021), particularly in the gig economy. Gig workers actively discuss their experiences on anonymous forums, e.g., Reddit, which have become a rich data source for researchers. Previous studies on rideshare worker concerns using Reddit, highlighting issues of privacy, scams, and support systems (Rosenblat and Stark, 2016; Sannon et al., 2022; Yao et al., 2021; Watkins, 2022; Ma et al., 2018), have been small-scale, limited by manual coding222the largest dataset analyzed about 2.6K posts (Watkins, 2022) using qualitative data analysis methods. We use our LLM-based framework to examine over 1 million comments from rideshare workers, significantly broadening the scope to amplify all voices.
3. QuaLLM Framework
@@
We introduce QuaLLM (Figure 1), a framework consisting of a prompting methodology and evaluation strategy for the analysis of text data from online forums.
3.1. LLM Prompting Methodology for Analysis of Online Forum Data
@@
Our approach begins with the collection of datasets from specified periods across any public threaded discussion forum, such as Reddit, Mastodon, and Facebook. We aim to extract and summarize the predominant concerns of workers within these communities through a four-step prompting process: generation, classification, aggregation, and prevalence.
3.1.1. Generation Prompt.
@@
The generation step involves the generation of concern summaries from online discussions relevant to a specific topic and linking each concern with a representative quote. The prompt is broken down into seven steps, following best prompt engineering practices from prior research and industry (OpenAI, 2023; Adams et al., 2023; Nori et al., 2023; Wang et al., 2023). The steps include:
  1. (1)
@@
  3. (3)
    selecting representative quotes333we argue that forcing the model to choose and output quotes is a form of Chain-of-Thought prompting (Wei et al., 2022) strategy and find that it helps prevent hallucination
  4. (4)
@@
To minimize API calls, we aggregate posts and comments into groups. For each group, we programmatically enhance the JSON output with the timestamp of the earliest post and assign a unique identifier, ensuring precise identification.
All these steps are necessary since direct prompting may cause hallucinations, redundancy, and inconsistent output. This prompting strategy is similar to “open coding,” where one derives new theories and concepts informed from the underlying data (Saldaña, 2021).
3.1.2. Classification Prompt.
We define a set of primary themes, typically around 4-5, based on prior research, parallel participant interviews, or other expert input, and the outputs of the generation prompt. We then prompt the LLM to classify each identified concern accordingly, with an additional “other” category for irrelevant concerns. This step mirrors “thematic analysis” (Braun and Clarke, 2012), which groups together a number of related codes, serving as a filtration mechanism.
3.1.3. Aggregation Prompt.
Aggregates concerns into secondary themes based on classifications, and identifies the most representative concerns (e.g., top 5) within each primary theme. We can manually select representative quotes for each secondary theme.
3.1.4. Prevalence Prompt.
Classifies concerns within primary themes as belonging to one of the secondary themes or an “other” category, facilitating quantitative analysis of concern prevalence by theme percentage.
3.2. LLM Outputs Evaluation Strategy
To validate the LLM outputs we propose a mix of human and computational evaluation methods based on the following metrics: factuality and completeness for generation, accuracy for classification and prevalence, and distinctness and coverage for aggregation prompts, respectively.
3.2.1. Generation Prompt:
@@
Factuality: Determines if the LLM’s output concerns reflect the source data. Evaluators respond to: “Is this candidate concern (and associated quote) output by the LLM present in the reference human generated list of transparency concerns? Answer Yes or No”. The factuality score is the proportion of accurate LLM-generated concerns, akin to precision.
Completeness: Evaluates the LLM’s ability to capture all relevant concerns, asking: “Is this reference human concern present in the candidate concern (and associated quote) output by the LLM? Answer Yes or No.” The completeness score denotes the proportion of human-identified concerns also recognized by the LLM, akin to recall.
3.2.2. Classification and Prevalence Prompts:
@@
3.2.3. Aggregation Prompt:
This prompt generates aggregated text given source text as input. We leverage a topic model to generate topics for the entire text and find those topics most similar to the subthemes. Human evaluation on a subset does not accurately capture the task, and instead we pivoted to a computational evaluation on the entire text using well established topic modeling techniques.
@@ 3.2.3. Aggregation Prompt:
Coverage (k): We measure the proportion of unique most-similar topics associated with each of the ‘n’ sub-themes that are also found among the most frequent ‘nk’ topics of the overall text. For example, if n=5 and k=2, we would look at the 5 sub-themes amongst the top 10 most frequent overall topics. A higher CR-k indicates that the sub-themes represent prevalent or significant themes in the text.
4. Case Study of Worker Concerns from Rideshare Subreddits
@@
We apply our framework to rideshare subreddit data (r/uberdrivers, r/lyftdrivers) using the GPT4-Turbo model on Microsoft Azure to identify worker concerns regarding AI and platform algorithms. See Appendices A, B, and C for our prompts, findings and costs incurred, respectively.
@@
Rideshare subreddits were chosen for their active 400K+ member community, presence of active discussions of worker concerns; which is of interest to regulators, accessible and comprehensive data, and relevance to HCI, Economics, and Communications researchers studying the harms caused to workers as a result of rideshare platforms’ algorithmic management (Zhang et al., 2022; Dubal, 2023; Rosenblat and Stark, 2016).
We now describe our dataset and summarize our findings.
4.1. Dataset and Experiments
@@
We obtained comprehensive 2019-2022 data from r/uberdrivers and r/lyftdrivers via https://the-eye.eu/redarcs/, focusing on this period to capture recent concerns, particularly relevant due to app changes and the unavailability of post-2022 data due to Reddit API changes [^4] Reddit API restrictions enforced in Feb’23; see https://www.redditinc.com/policies/data-api-terms. After post pre-processing to remove short samples, our dataset comprised 65,377 submissions (47,1064710647,10647 , 106 from Uber, 18,2711827118,27118 , 271 from Lyft) and 1,392,77613927761,392,7761 , 392 , 776 associated comments (1,054,03010540301,054,0301 , 054 , 030 from Uber, 338,746338746338,746338 , 746 from Lyft) [^5] On Reddit, a submission is a post, and a comment is a response to a post or another comment; we concatenate and use the text from posts and all associated comments. We then grouped submissions and comments into sets of five for each LLM API call (See Appendix A for the prompts). We encountered an 11% error rate due to throttling and content issues. Finally we obtained 58,728 concerns which we categorized into four themes: Enhancing Transparency and Explainability (24,721 concerns, 42%), Predictability and Worker Agency (12,728 concerns, 22%), Better Safety and More Time (6,144 concerns, 10.5%), and Ensuring Fairness and Non-Discrimination (4,280 concerns, 7%); the remaining 18.5% were categorized as “Other”. We then aggregated these findings to highlight the top five concerns for each theme and calculated their prevalence across the dataset.
4.2. Findings
@@
We briefly summarize our findings below. We also provide tables of the aggregated, ranked and prevalence quantified themes in Appendix B.
@@
(i) Enhancing Transparency and Explainability: Driver concerns about opaque fare calculations, unclear incentives, and uncertain criteria for earning, surge pricing, and cancellation (Table 1)
@@
4.3. Evaluation
4.3.1. Generation Prompt.

### v2

@@ Abstract
Online discussion forums provide crucial data to understand the concerns of a wide range of real-world communities. However, the typical qualitative and quantitative methodologies used to analyze those data, such as thematic analysis and topic modeling, are infeasible to scale or require significant human effort to translate outputs to human readable forms. This study introduces QuaLLM, a novel LLM-based framework to analyze and extract quantitative insights from text data on online forums. The framework consists of a novel prompting and human evaluation methodology. We applied this framework to analyze over one million comments from two of Reddit’s rideshare worker communities, marking the largest study of its type. We uncover significant worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls about worker insights. In short, our work sets a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums.
@@
QuaLLM: An LLM-based Framework to Extract Quantitative Insights from Online Forums
Varun Nagaraj Rao Princeton University / CITP varunrao@princeton.edu                        Eesha Agarwal Princeton University eagarwal@alumni.princeton.edu
Samantha Dalal University of Colorado, Boulder samantha.dalal@colorado.edu                        Dana Calacci Penn State University dcalacci@psu.edu                        Andrés Monroy-Hernández Princeton University / CITP andresmh@princeton.edu
1 Introduction
@@
Text data from online forums, such as Reddit, has been used to study issues affecting diverse communities Proferes et al. (2021). Researchers have leveraged this data across various domains, including healthcare Zamanifard and Robb (2023); Xu et al. (2023); Ho et al. (2017); Paxman (2021), political discourse Papakyriakopoulos et al. (2023); De Francisci Morales et al. (2021); Marchal (2020), education Madsen and Madsen (2022); Park et al. (2018), and gig work [^1] A market characterized by short-term and freelance work like food delivery and ridesharing Rosenblat and Stark (2016); Sannon et al. (2022); Yao et al. (2021); Watkins (2022); Ma et al. (2018), to highlight critical concerns and trends.
Figure 1: QuaLLM transforms large-scale unstructured text-based online forum discussions on platforms like Reddit and Facebook into a structured survey-style format, identifying top-level themes associated with prevalence ranked sub-themes (by frequency of occurrence) and representative quotes.
Figure 2: QuaLLM’s multiphase prompting attempts to evoke the steps that human analysts might perform.
@@
Studies analyzing this text data for social phenomena often use interpretive qualitative methodologies Gichuru (2017) like coding Saldaña (2021) and thematic analysis Braun and Clarke (2012), to extract themes. Although detailed, these methodologies demand considerable human effort and interpretation. Such detailed interpretations may not always be necessary for identifying broad trends across large online communities. Alternatively, positivist quantitative Park et al. (2020) techniques like LDA Abdul et al. (2018) and BERT derivatives for topic modeling distill topics, which humans can interpret and convert into themes  [^2] We distinguish between topics/concepts and themes; a topic/concept is a concrete and specific subject matter of a text often expressed as a single word or a short phrase, while a theme is an interpretive and abstract insight conveyed by the text, typically articulated as a complete sentence., offering greater scalability. However, they require careful hyperparameter selection, operate on limited context length, and require significant human effort to translate outputs to human readable forms.
@@
Recent advancements in LLMs and prompt engineering Shah et al. (2023); Nori et al. (2023); OpenAI (2023); Deng et al. (2023) offer an alternative for analyzing extensive text data from online forums, addressing limitations of scale and context. Specifically, LLMs have been shown to perform well on generation, classification, and ranking tasks Zhao et al. (2023), and thus have the potential to assist researchers in text data analysis. However, the lack of an effective methodology for extracting themes from online communities limits their use. We address this with the following contributions:
@@
    We introduce QuaLLM, an LLM-based framework consisting of a novel prompting and human evaluation methodology for the thematic analysis and extraction of quantitative insights from online forums’ text data (See Figure 1).
@@
    We apply our framework to a case study on Reddit’s rideshare communities, analyzing over one million comments—the largest study of its kind—to identify worker concerns regarding AI and algorithmic platform decisions, responding to regulatory calls White House OSTP (2023).
@@
Taken together, our work establishes a new precedent for AI-assisted quantitative data analysis to surface concerns from online forums.
2 Related Work
@@
LLM Prompting and Evaluation.   LLMs have been used for generation of topics and concepts Shah et al. (2023); Zhao et al. (2023) [^3] The novelty of our work lies in surfacing themes (e.g., “facing fear leads to growth”) rather than topics or concepts (e.g., “fear,” “love,” “bias of LLMs”), classification Deng et al. (2023), and ranking Qin et al. (2023) tasks via advances in prompt engineering Nori et al. (2023); OpenAI (2023). Popular prompting methods include: In-context learning Brown et al. (2020), which teaches LLMs new tasks with few labeled examples, Chain of Thought Wei et al. (2022), which breaks down complex problems through stepwise reasoning, and EnsemblingWang et al. (2022) which aggregates outputs of multiple model runs to obtain a consensus output. Despite these advances, accurately evaluating LLM outputs is challenging, with human evaluation generally preferred Amer (2023). Building on prior work, we introduce a framework consisting of a novel prompting and human evaluation methodology for the quantitative analysis of text data from online forums.
@@
Gig Work and Reddit Analysis.   Regulatory bodies White House OSTP (2023) have sought public insights on the impact of AI and algorithmic management by employers Kantor and Sundaram (2022); Bernhardt et al. (2021), particularly in the gig economy. Gig workers actively discuss their experiences on anonymous forums, e.g., Reddit, which have become a rich research data source. Previous studies on rideshare worker concerns using Reddit, highlighting issues of privacy, scams, and support systems Rosenblat and Stark (2016); Sannon et al. (2022); Yao et al. (2021); Watkins (2022); Ma et al. (2018), have been small-scale, limited by manual coding444the largest dataset analyzed about 2.6K posts Watkins (2022) using qualitative data analysis methodologies. We use QuaLLM to examine over 1 million comments from rideshare workers, significantly broadening the scope to amplify all voices and also to translate this academic research into policy-relevant insights Nagaraj Rao et al. (2024a, b, 2025).
3 QuaLLM Framework
@@
We introduce QuaLLM (Figure 2), a framework consisting of a prompting method and evaluation strategy for the thematic analysis of text data from online forums.
3.1 LLM Prompting Method for Analysis of Online Forum Data
@@
Our approach begins with collecting datasets from specified periods across any public threaded discussion forum, such as Reddit and Facebook. We aim to extract and summarize the predominant concerns of workers within these communities through a four-step prompting process: generation, classification, aggregation, and prevalence.
@@
Generation Prompt.   The generation step involves the generation of concern summaries from online discussions relevant to a specific topic and linking each concern with a representative quote. The prompt is broken down into seven steps, following best prompt engineering practices from prior research and industry OpenAI (2023); Adams et al. (2023); Nori et al. (2023); Wang et al. (2023). The steps include:
  1. 1.
@@
  3. 3.
    selecting representative quotes555we argue that forcing the model to choose and output quotes is a form of Chain-of-Thought prompting Wei et al. (2022) strategy and find that it helps prevent hallucination
  4. 4.
@@
To minimize API calls, we aggregate posts and comments into groups. For each group, we programmatically enhance the JSON output with the timestamp of the earliest post and assign a unique identifier. All these steps are necessary since direct prompting may cause hallucinations, redundancy, and inconsistent output. This prompting strategy is similar to “open coding,” where one derives new theories and concepts informed from the underlying data Saldaña (2021).
Classification Prompt.   We define a set of top-level themes, typically around 4-5, based on prior research, parallel participant interviews, or other expert input, and the outputs of the generation prompt. We then prompt the LLM to classify each identified concern accordingly, with an additional “other” category for irrelevant concerns. This step mirrors “thematic analysis” Braun and Clarke (2012), which groups together a number of related codes, serving as a filtration mechanism.
Aggregation Prompt.   Aggregates concerns into sub-themes based on classifications and identifies representative concerns (e.g., top 5) within each top-level theme. We manually select representative quotes for each sub-theme.
Prevalence Prompt.   Classifies concerns within top-level themes as belonging to one of the sub-themes or an “other” category, facilitating quantitative analysis of concern prevalence by theme percentage.
3.2 LLM Outputs Evaluation Strategy
To validate the LLM outputs, we propose a mix of human and computational evaluation methods based on the following metrics: factuality and completeness for generation, accuracy for classification and prevalence, and distinctness and coverage for aggregation prompts, respectively.
3.2.1 Generation Prompt:
@@
Factuality: Determines if the LLM’s output concerns reflect the source data. Evaluators respond to: “Is this candidate concern (and quote) output by the LLM present in the reference human generated list of transparency concerns? Answer Yes or No”. The factuality score is the proportion of accurate LLM-generated concerns akin to precision.
Completeness: Evaluates the LLM’s ability to capture all relevant concerns, asking: “Is this reference human concern present in the candidate concern (and associated quote) output by the LLM? Answer Yes or No.” The completeness score denotes the proportion of human-identified concerns recognized by the LLM, akin to recall.
3.2.2 Classification and Prevalence Prompts:
@@
3.2.3 Aggregation Prompt:
This prompt generates aggregated text given source text as input. We leverage a topic model to generate topics for the entire text and find those topics most similar to the subthemes. Human evaluation on a subset does not accurately capture the task, and instead, we pivoted to a computational evaluation of the entire text using well-established topic modeling techniques.
@@ 3.2.3 Aggregation Prompt:
Coverage (k): We measure the proportion of unique most-similar topics associated with each of the ‘n’ sub-themes that are also found among the most frequent ‘nk’ topics of the overall text. For example, if n=5 and k=2, we would look at the 5 sub-themes amongst the top 10 most frequent overall topics. A higher CR-k implies the sub-themes represent prevalent or significant themes in the text.
4 Case Study of Worker Concerns from Rideshare Subreddits
@@
We apply our framework to rideshare subreddit data (r/uberdrivers, r/lyftdrivers) using the GPT4-Turbo model on Microsoft Azure to identify worker concerns regarding AI and platform algorithms. See Appendices A, B, and D for our prompts, findings, and costs incurred, respectively.
@@
We chose rideshare subreddits for their active 400K+ member community, presence of active discussions of worker concerns, which is of interest to regulators, accessible and comprehensive data, and relevance to Human-Computer Interaction, Economics, and Communications researchers studying the harms caused to workers as a result of rideshare platforms’ algorithmic management Nagaraj Rao et al. (2025); Zhang et al. (2022); Dubal (2023); Rosenblat and Stark (2016). We now describe our dataset and summarize our findings.
4.1 Dataset and Experiments
@@
We obtained comprehensive 2019-2022 data from r/uberdrivers and r/lyftdrivers via https://the-eye.eu/redarcs/, focusing on this period to capture recent concerns, particularly relevant due to app changes and the unavailability of post-2022 data due to Reddit API changes [^6] Reddit API restrictions enforced in Feb’23; see https://www.redditinc.com/policies/data-api-terms. After processing to remove short samples, our dataset [^7] On Reddit, a submission is a post, and a comment is a response to a post or another comment. We concatenate and use the text from posts and all associated comments comprised 65,377 submissions (47,1064710647,10647 , 106 from Uber, 18,2711827118,27118 , 271 from Lyft) and 1,392,77613927761,392,7761 , 392 , 776 associated comments (1,054,03010540301,054,0301 , 054 , 030 from Uber, 338,746338746338,746338 , 746 from Lyft). We then grouped submissions and comments into sets of five for each LLM API call (See Appendix A for the prompts). We encountered an 11% error rate due to throttling and content issues. Finally, we obtained 58,728 concerns which we categorized into four themes: Enhancing Transparency and Explainability (24,721 concerns, 42%), Predictability and Worker Agency (12,728 concerns, 22%), Better Safety and More Time (6,144 concerns, 10.5%), and Ensuring Fairness and Non-Discrimination (4,280 concerns, 7%); the remaining 18.5% were categorized as “Other”. We then aggregated these findings to highlight the top five concerns for each theme and calculated their prevalence across the dataset.
4.2 Findings
@@
We briefly summarize our findings below. We also provide tables of the aggregated, ranked, and prevalence-quantified themes in Appendix B.
@@
(i) Enhancing Transparency and Explainability: Drivers are concerned about opaque fare calculations, unclear incentives, and uncertain criteria for earnings, surge pricing, and cancellations (Table 2)
@@
4.3 Evaluation
Evaluation Metrics Factuality Completeness Distinctness Coverage (k) Accuracy
Analysis phase
evaluated with
metric
Generation Generation Aggregation Aggregation
Classification,
Prevalence
Reference data
generation
method
Theme extraction
by human
(2 researchers)
Theme extraction
by human
(2 researchers)
Topic extraction using
topic models
Topic extraction using
topic models
Labeling
by humans
(2 researchers)
Comparison b/w
reference data and
LLM output
Manually
(2 researchers)
Manually
(2 researchers)
Automatically
(Python Program)
Automatically
(Python Program)
Automatically
(Python Program)
Limitation of
reference data
generation method
Costly Costly
Need for interpretation
of topic model outputs
Need for interpretation
of topic model outputs
Costly
Evaluation
sample size
125 posts
2,511 comments
125 posts
2,511 comments
47,873 concerns 47,873 concerns 100 concerns
Actual data size
65,377 posts
1,392,776 comments
65,377 posts
1,392,776 comments
47,873 concerns 47,873 concerns
58,728 concerns
47,873 concerns
Results 0.55* 0.78* 0.80*
0.95* (k=1)
1.00* (k=2)
0.74* (classification)
0.82* (prevalence)
Table 1: Metrics Chosen to Evaluate LLM-Powered Analysis Phases and Results. Metric values belong to the closed interval [0,1], with higher values indicative of better performance. * values are statistically significantly different from chance measured via a binomial test (p-value <0.05).
4.3.1 Generation Prompt.